OpenStudy (anonymous):

HCN(aq)+SO4-2(aq)HSO4-(aq)+CN-(aq) What is the Bronsted-Lowry acid in this equation? What is the Bronsted-Lowry Base in this equation? What is the Conjugate acid in this equation? What is the Conjugate base in this equation?

OpenStudy (aaronq):

Bronsted acids donate H+ Bronsted bases accept H+ CB=conjugate base CA=conjugate acid acid base CB CA HA + B- <-> A- + BH

OpenStudy (anonymous):

Bronsted lowry acid=HCN(aq) Bronstes lowry base =SO4-2 CB=HSO4- CA=CN- is this correct?

OpenStudy (aaronq):

yep, this is right